A biquadratic equation f(x) is having zeroes -2,3,5, and 7..and f(0) is 1. so the remainder of f(x) when divided by (x-1) is...please help with explanation..

Grade:12

2 Answers

Vikas TU
3 years ago
Der student 
It is not a part of JEE syllabus 
Please attach an image , we will happy to help you 
Good Luck 
Cheers 
Aditya Gupta
3 years ago
 
dear student, biquad polynomial means it has degree 4.
also, since its roots are given to be – 2, 3, 5, 7, it means that we can write f(x) as follows:
f(x)= A(x – ( – 2))(x – 3)(x – 5)(x – 7)
= A(x+2)(x – 3)(x – 5)(x – 7)
put x=0 both sides
f(0)= A(2)(-3)(-5)(-7)
or A= – 1/210
so, the remainder when f(x) is divided by x – 1, would be f(1) (using remainder theorem)
so, f(1)= A(1+2)(1 – 3)(1 – 5)(1 – 7)
= – (1/210)(3)(-2)(-4)(-6)
24/35
option C
KINDLY APPROVE :))